OpenStudy (anonymous):

how to solve x+4y=-9 4x-3y=2 by using elimination method

OpenStudy (shamil98):

x + 4y = -9 4x - 3y = 2 you want to get rid of x first because it is the easiest method and requires less work so you multiply the top equation by -4 -4(x+4y=-9) -4x -16y = 36 4x -3y = 2 0 - 19y = 38 -19y= 38 y = -2 and solve for x from there.

OpenStudy (anonymous):

x + 4y = -9 x + 4(-2) = -9 x -8 = - 9 x = -9 + 8 x = -1, right?

OpenStudy (shamil98):

Yes. Good job.